Status Report

ISS On-Orbit Status 26 Jul 2002

By SpaceRef Editor
July 26, 2002
Filed under , ,

All ISS systems continue to function nominally, except as noted previously
or below.

The crew last night acknowledged that yesterday’s emergency/fire drill OBT
(on-board training) was beneficial. Flight Control: “Staying well trained
is the best way to ensure you never need it.”

Today, FE-1 Peggy Whitson worked on installation of the alignment guides in
the ARIS (active rack isolation system) of EXPRESS Rack 2 (ER2). [Because
of the difficulty each crew has had installing the ARIS alignment guides,
the ground uplinked some questions regarding the installation. Crew
feedback will help the engineers better understand what can be done
operationally to make installation easier, which then will help to make the
best determination of what hardware modification or redesign should be

Because of the continuing troubleshooting of the SUBSA (solidification using
a baffle in sealed ampoules) experiment, POC (Payload Operations Center)
decided to abort SUBSA activities for today and 7/29 for the time being.
SUBSA team is performing ground tests on the Eurotherm controller and will
solidify a recovery plan.

FE-2 Sergei Treschev performed the periodic task of replacing dust filters
in the DC-1 docking compartment, while CDR Valery Korzun replaced four dust
filters (PF1-4) in the SM air circulation system.

Peggy Whitson’s schedule today include the repeat of the photo documenting
of the ADVASC (advanced astroculture) growth chamber. From the photos taken
on 7/23, the ADVASC team could not conclusively identify the developmental
stage of the flowers and pod filling. When Peggy repeated the plant photo
activity, she used the manual focus mode on the camera to take pictures of
individual pods and flowers.

Using toys, the crew playfully demonstrated (and video-recorded) physical
principles in zero-G, as part of the EPO (Educational Payload Operations)
program. [They were requested to repack the toys after completing
operations in the original packing configuration to avoid damage (some items
are fragile and may be damaged during re-entry, because the Principal
Investigator (PI) wants to include some of the toys in a museum collection.]

The ground uplinked additional explanations of the on-going TVIS (treadmill
with vibration isolation and stabilization) troubleshooting, i.e., tracking
down an anomalous noise heard by the crew in the treadmill on 7/24. [Based
on crew answers to subsequent questions, the engineering team agreed that
until symptoms were understood more fully, no further on-orbit TVIS
troubleshooting should be performed. This is because there is the potential
of making the problem worse by re-energizing the TVIS and that there was
some risk of completely losing TVIS belt operation, including passive mode
exercise. Further on-orbit troubleshooting was put on hold until the ground
hardware could be prepared and tested. On 7/27 or later, a conference will
be held with the crew on goals/objectives of a newly prepared procedure.

FE-2 Treschev performed the daily routine tasks of SOSh life support systems
maintenance and IMS (inventory management system) preparation, while CDR
Korzun did the daily payload status check in the Lab.

The Russian TsVM (Central Computer system) lost its second subset. Of the
three redundant “lanes”, only lane 2 is operating. The TVM (Terminal
Computer system) continues to run on lanes 1 and 3, with one subset dropped
out of the set.

The solar Beta angle of the ISS orbit is now at 70.6 deg and continues to
increase toward its yearly maximum of 72 deg.

SpaceRef staff editor.